### v1

@@ 5 Results
Figure 1: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the No Admit, Very Short Stay, Short Stay, & Long Stay class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm 2\sigma$) of the feature’s impact on the prediction. The plot suggests that NumTransfers, NumNotes, AdmitProcedure are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the No Admit class; NumTransfers, NumNotes, NumDiagnosis are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Very Short Stay class; NumNotes, NumTransfers, AdmitProcedure are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Short Stay class; NumNotes, NumTransfers, NumDiagnositcs are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Long Stay class.
(a) NumTransfers
(b) NumNotes
(c) AdmitProcedure
Figure 2: The distribution of SHAP values for the four most contributing features to the No Admit class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) NumTransfers
(b) NumNotes
(c) NumDiagnosis
Figure 3: The distribution of SHAP values for the four most contributing features to the Very Short Stay class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) NumNotes
(b) NumTransfers
(c) AdmitProcedure
@@
(a) NumNotes
(b) NumTransfers
(c) NumDiagnosis
@@
The Medical Information Mart for Intensive Care III (MIMIC-III) [78] is a substantial clinical database comprising detailed health-related data from over 40,000 adult patients admitted to critical care units at a tertiary care hospital between 2001 and 2012. The dataset under consideration is extensive in nature, encompassing a wide range of information pertinent to the subject. This includes demographic data, vital signs, laboratory test results, medications, procedures, clinical notes, imaging reports, and hospital length of stay. The features ’hadm_id’, ’LOSdays’, ’religion’, ’marital_status’, ’ethnicity’ were removed for ethical and privacy reasons and to replicate the real-world scenarios. We utilised this dataset to classify the datapoints into length of stay (LOS). We then, grouped the classes into No Admission, Very Short Stay, Short Stay, and Long Stay categories. No Admission applies to patients who are assessed but not admitted to the hospital. Very Short Stay refers to hospitalizations lasting less than three days, while Short Stay covers stays from three to seven days. Long Stay describes hospitalizations exceeding seven days [79]. The feature descriptions for the features we have discussed are as follows, NumTransfers represents the number of times a patient is transferred within the hospital during their stay. These transfers may occur between different units, such as from the emergency department to the ICU, or between wards. They reflect patient movement within the hospital.NumNotes is the count of clinical notes recorded for a patient during their hospital admission or ICU stay. These notes include documentation from physicians, nurses, and other healthcare providers, capturing clinical observations, treatments, and progress. Admit Procedures is number or list of procedures performed around the time of hospital admission. These procedures could be diagnostic or therapeutic interventions documented in the procedure coding system and reflect initial clinical care. NumDiagnosis is the number of unique diagnoses assigned to the patient during their hospital or ICU stay. Diagnoses are typically coded using International Classification of Diseases (ICD) codes, which provide an overview of the patient’s clinical conditions. gender is the administrative gender identity of the patient, as recorded in hospital records. In MIMIC, this is usually "M" (male) or "F" (female), though other categories may be included depending on the data source. ExpiredHospital is a binary flag indicating whether the patient died during the hospital stay (1 for yes and 0 for no). This captures in-hospital mortality as a key outcome variable.
The RF Classifier was trained with 100 numbers of estimators along with the default parameters following an 80:20 train test stratified split. The reported F1 score on the dataset was 89.0%. The implementation of the proposed framework on the test data set was undertaken to assess the epistemic uncertainty inherent in predictions pertaining to unseen data. The number of samples, denoted by the parameter "number of samples," was set to 100 sub-models or tree subsets. This was done to simulate 100 posterior samples from the model space. The parameter designated as $\alpha$ is employed to modulate the degree to which sub-trees are explored. In the context of model-based search algorithms, the parameter $\alpha$ serves to determine whether the user’s objective is to identify balanced trees, trees with optimal performance, or trees characterized by uniform uncertainty estimation. For this experiment, the value of $\alpha$ is selected to be less than one, with the objective of identifying the trees with the highest performance. For each of the classes, the absolute SHAP values are calculated (see Figure:  1).
[…]
Furthermore, we examined the top three features, their associated epistemic uncertainty, sign stability, and SHAP value distribution, as well as the mean SHAP and 95% SHAP confidence interval. For the No Admit or No Admission category, the epistemic uncertainties for the features NumTransfers, NumNotes, and AdmitProcedures are 0.063, 0.061, and 0.045, respectively (Figure: 2). The sign stability for these features is 99.0%, 72.0%, and 96.0%, respectively. These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. Explanation entropy for the features NumTransfers is considerably low as well, indicated by a uniform distribution of SHAP values, signaling high information certainty about the feature’s impact. The feature AdmitProcedure has high entropy, indicated by skewed SHAP value distributions. NumNotes have sign stability of 72.0%, indicating interpretive inconsistency despite low epistemic uncertainty. For the Very Short Stay category, the epistemic uncertainties for the features NumTransfers, NumNotes, and NumDiagnosis are 0.063, 0.055, and 0.053, respectively. The sign stabilities for these features are 100.0%, 99.0%, and 96.0%, respectively (Figure 3). The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The sign stabilities are very high, indicating interpretive consistency with low epistemic uncertainty. For the Short Stay category, the epistemic uncertainties for the features NumNotes, NumTransfers, and AdmitProcedures are 0.060, 0.038, and 0.048, respectively. The sign stabilities for these features are 84.0%, 85.0%, and 89.0%, respectively (Figure: 4). The standard deviation of epistemic uncertainties is low in the top three features, suggesting consistent SHAP values across the ensemble and reflecting certainty about the features’ influence. Explanation entropy for the features NumTransfer, AdmitProcedures, & NumNotes is considerably high as well, indicated by a non-uniform distribution of SHAP values, signaling high information certainty about the feature’s impact. The sign stabilities are mostly stable, indicating interpretive inconsistency despite low epistemic uncertainty. For the Long Stay category, the epistemic uncertainties for the features NumNotes, NumTransfers, and NumDiagnosis are 0.041, 0.033, and 0.030, respectively. The sign stabilities for these features are 100.0%, 99.0%, and 90.0%, respectively (Figure:  5). The standard deviation of epistemic uncertainties is low in the top three features, suggesting consistent SHAP values across the ensemble and reflecting certainty about the features’ influence. However, a skewed distribution of SHAP values signals low information certainty about the feature’s impact. The features have sign stability, indicating interpretive consistency with low epistemic uncertainty.
Ovarian Cancer Dataset
(a) No Risk
(b) Low Risk
(c) Medium Risk
(d) High Risk
@@
Figure 6: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the No Risk, Low Risk, Medium Risk, & High Risk class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction. The plot suggests that Smoking, PreviousTreatment, & HormoneTherapy are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the No Risk class; Symptom, PreviousTreatment, & EnhancementPattern are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Low Risk class; Smoking, PreviousTreatment, & HormoneTherapy, are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Medium Risk class; Smoking, HormoneTherapy, & PreviousTreatment are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the High Risk class.
(a) Smoking
(b) PreviousTreatment
(c) HormoneTherapy
@@
The Ovarian Cancer Dataset [80] contains 200,100 patient records collected hourly between January 2019 and December 2024. This highly detailed longitudinal dataset is useful for monitoring ovarian cancer risk and progression. It is designed to support prognostic modeling and progression risk assessment in ovarian cancer patients. This dataset contains 200,100 data points and 34 features. It was used to categorize ovarian cancer in females into risk categories. For each data point, the associated feature, Risk Label, has four classes: No Risk, Low Risk, Medium Risk, and High Risk. No Risk coressponds to no evidence or indication of risk. Low Risk corresponds to minimal probability of an adverse outcome or malignancy. Medium Risk coressponds to moderate chance of risk; requires monitoring or further evaluation. High Risk indicates a high probability of an adverse outcome or malignancy and likely warrants intervention. The description of the features we have discussed in the results are as mentioned further. Symptom coressponds to clinical signs or patient-reported symptoms that are associated with the progression or risk of ovarian cancer and are used to characterize the disease. PreviousTreatment coressponds to information on any medical therapies or interventions the patient underwent before the current assessment, such as chemotherapy, surgery, or radiation therapy. EnhancementPattern is an imaging-derived feature that describes the contrast enhancement patterns observed in diagnostic scans. Smoking is the patient’s smoking history or status is a known risk factor impacting ovarian cancer progression and overall health. HormoneTherapy/HormoneTreatment are the records of hormonal treatments received by the patient, including exogenous hormone administration, which may influence cancer risk or progression. SocioeconomicStatus is a categorical or continuous measure reflecting the patient’s social and economic circumstances, which can affect access to healthcare and outcomes. SNP_Status is a genetic feature indicating the presence or absence of specific single nucleotide polymorphisms (SNPs) related to ovarian cancer susceptibility or progression.
The dataset was highly imbalanced, with the No Risk class having 119,965 data points, the Low Risk class having 40,092 data points, the Medium Risk class having 30,068 data points, and the High Risk class having 9,975 data points. We implemented oversampling to create a class balance using the hybrid resampling technique SMOTETomek [81], which combines oversampling using SMOTE [82] and undersampling using TOMEK link removal. This technique removes noisy and borderline instances after oversampling. The resulting dataset was then used to create an 80:20 training-testing split with stratified sampling. We trained an RF classifier with number of estimators equal to 500 and default parameters on the training set, achieving an F1 score of 58.0% on the test set. Our main focus with this dataset was studying uncertainty and analyzing it when the model cannot learn complex patterns from the dataset. We implemented our framework on the test data to analyze epistemic uncertainty in predictions on unseen data. For the experiment on this dataset, the 500 sub-models or tree subsets parameter for the number of samples was set to simulate 500 posterior samples from the model space. We chose the $\alpha$ parameter to be 1.0 to perform balanced sampling of the sub-ensembles. We calculated the absolute SHAP values for each risk label. For the No Risk class, the top three most important features were identified to be Smoking, PreviousTreatment, & HormoneTherapy(Figure 6). For the Low Risk class, the top three most important features were identified to be Symptom, PreviousTreatment, & EnhancementPattern. For the MediumRisk class the top three most important features were identified to be Smoking, HormoneTherapy, & PreviousTreatment. For the High Risk class the top three most important features were identified to be Smoking, HormoneTherapy, & PreviousTreatment. The results from the absolute mean SHAP values chart (Figure 6) indicate that SHAP feature importance varies considerably among the underlying models. This suggests considerable variability across different model subsets. While the model consistently relies on features such as Symptom, MenstrualHistory, SocioeconomicStatus, HormoneTreatment, etc. , it does so with substantial epistemic ambiguity regarding their precise impact magnitude. In contrast, features such as SNP_Status exhibit low average SHAP values and narrow uncertainty narrow violin plots, reflecting low importance and high confidence in their negligible contribution. This hints at stable but marginal roles in predicting the different classes. This feature has consistent performance across all classes. One major difference observed is in terms of positively or negatively influencing predictions.
We also studied the top three features of this dataset and their associated epistemic uncertainty, sign stability, and SHAP value distribution, as well as the mean SHAP and 95% SHAP confidence interval. For the No Risk category, the epistemic uncertainties for the features No Risk category the epistemic uncertainties for the features Smoking, PreviousTreatment, & HormoneTherapy are 0.050, 0.078, and 0.057, respectively (Figure 7). The sign stabilities for these features are 100.0%, 94.6%, and 100.0%, respectively. These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP distribution is non-uniform and skewed for the Smoking and HormoneTherapy features, signaling high information uncertainty about the feature’s impact. These features have high sign stability, indicating interpretive consistency and low epistemic uncertainty. For the Low Risk category, the epistemic uncertainties for the features Symptom, PreviousTreatment, & EnhancementPattern are 0.016, 0.063, and 0.020, respectively. The sign stabilities for these features are 98.4%, 71.6%, and 86.8%, respectively (Figure: 8). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP has a non-uniform distribution for the Symptom and EnhancementPattern features, signaling high information uncertainty about the feature’s impact. The features PreviousTreatment and EnhancementPattern have a sign stability of 71.6% and 86.8%, respectively, indicating not-very-high interpretive consistency despite low epistemic uncertainty. The SHAP distribution exhibits flatness, indicating disagreement across the sub-ensembles. For the Medium Risk category, the epistemic uncertainties for the features Smoking, HormoneTherapy, & PreviousTreatment are 0.050, 0.015, and 0.051, respectively. The sign stabilities for these features are 92.0%, 97.2%, and 82.8%, respectively (see Figure: 9). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP has a non-uniform distribution for the Smoking and HormoneTheraphy features, signaling high information uncertainty about the feature’s impact. The featurePreviousTreatment have a sign stability of 82.8% indicating not-very-high interpretive consistency despite low epistemic uncertainty. The SHAP distribution exhibits flatness, indicating disagreement across the sub-ensembles. For the High Risk category, the epistemic uncertainties for the features Smoking, PreviousTreatment, & HormoneTherapy are 0.026, 0.036, and 0.040, respectively. The sign stabilities for these features are 100%, 75.4%, and 96.6%, respectively (see Figure: 10). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP distribution is highly skewed for all the features, signaling high information uncertainty about the feature’s impact, along with a relatively dispersed distribution. This reinforces the model’s uncertainty about precise attribution. The features have a very high sign stability of 75.4%, indicating high interpretive consistency with low epistemic uncertainty.
SEER Breast Cancer Dataset
(a) Deceased Patients
(b) Surviving Patients
@@
Figure 11: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the Deceased & Alive class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction. The plot suggests that Age, Regional_Node_Examined, & Regional_Node_Positive are the most impactful feature, significantly contributing to the high probability of the prediction’s shift towards the both Deceased & Alive class, although in different directions.
(a) Age
(b) Regional_Node_Examined
(c) Regional_Node_Positive
@@
Figure 12: The distribution of SHAP values for the four most contributing features to the Alive class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) Age
(b) Regional_Node_Examined
(c) Regional_Node_Positive
@@
The SEER Breast Cancer Dataset [83] is a comprehensive cancer registry that provides extensive information on breast cancer cases, including patient demographics, tumor characteristics, treatment details, and survival outcomes. The utilization of this method is prevalent in the analysis of treatment outcomes, as it captures comprehensive, population-based longitudinal data. This capability enables researchers and clinicians to assess the efficacy of diverse interventions across a range of patient populations and clinical settings. For instance, advanced predictive models developed using SEER data apply machine learning to predict individual patient survival and treatment response, enabling personalized treatment decisions and optimizing treatment strategies to extend survival while minimizing adverse effects. This capability is of particular significance in complex cases, such as metastatic breast cancer, where the guidance provided by clinical trials is limited and SEER-based models facilitate decision-making by simulating the outcomes of various treatment options that are tailored to the unique characteristics of each patient and tumor [84, 85]. The dataset contains 4,024 data points, 16 features, and associated survival outcomes, specifically the Alive and Dead status, for female patients diagnosed with breast cancer.
The Alive class indicates the patients who were alive at the final follow-up or censoring time. Dead class indicates patients who died from breast cancer or any other cause during the observation period. The features under discussion are described further text. Age: The age of the patient at the time of the initial breast cancer diagnosis. This influences prognosis and treatment choice. Regional_Node_Positive: The number of regional lymph nodes confirmed positive for cancer, indicating spread. Grade_Poorly differentiated (Grade III): Tumor cells are highly abnormal and grow/spread aggressively. Grade_Moderately differentiated (Grade II): Tumor cells are highly abnormal and grow/spread aggressively. Regional_Node_Examined indicating that the total number of regional lymph nodes examined for cancer involvement. Progesterone_Status_Negative indicating that the tumor cells lack progesterone receptors, which may affect response to hormone therapy. T_Stage_T4 corresponding that the tumor has invaded the chest wall and/or skin (advanced tumor size/invasion stage). N_Stage_N1 denoting that the cancer has spread to one to three axillary lymph nodes. N_Stage_N2 indicating that the cancer has spread to four to nine axillary lymph nodes, or the nodes are fixed or matted.N_Stage_N3 indicating that the cancer has spread to 10 or more axillary nodes, as well as to the infraclavicular or internal mammary nodes. 6th_Stage_IIA is the AJCC 6th edition Stage IIA indicating moderately advanced local disease.6th_Stage_IIB is the AJCC 6th edition Stage IIB indicating larger tumor size and/or more node involvement.6th_Stage_IIIA is the AJCC 6th edition Stage IIIA indicating advanced local spread to several regional nodes.6th_Stage_IIIB is the AJCC 6th edition Stage IIIB indicating tumor involves chest wall or skin and may involve nodes.6th_Stage_IIIC is the AJCC 6th edition Stage IIIC indicating extensive lymph node involvement near collarbone or breastbone. The dataset was then employed for the purpose of survival classification to study the treatment outcome. The 80:20 stratified train-test split was implemented. Subsequently, an RF (random forest) classifier with a number of estimators equal to 300 along with default parameters, was trained. The model was evaluated on the test set, and the resultant F1 score was 79.8%. The implementation of the proposed framework on the test data set was undertaken to assess the epistemic uncertainty inherent in predictions pertaining to test data. The value of the $\alpha$ parameter was set to 0.5 to perform the balanced exploration of the trees.
The absolute mean SHAP values chart (Figure: 11) shows that the SHAP feature importance varies considerably among the top contributing features in the underlying models. This suggests considerable variability across different model sub-ensembles. While the model consistently relies on features such as Age, Regional_Node_Positive, Grade_Poorly differentiated: Grade III, Grade_Moderately differentiated: Grade II, Regional_Node_Examined, Progesterone_Status_Negative, the model exhibits substantial epistemic ambiguity regarding their precise impact magnitude. In contrast, features such as T_Stage_T4, N_Stage_N1, N_Stage_N2, N_Stage_N3, 6th_Stage_IIA, 6th_Stage_IIB, 6th_Stage_IIIA, 6th_Stage_IIIB, 6th_Stage_IIIC exhibit low average SHAP values and narrow uncertainty violin plots. This reflects both low importance and high confidence in their negligible contribution, hinting at stable but marginal roles in predicting the different classes. One major difference observed is in terms of positively or negatively influencing predictions. For each survival status, we calculate the absolute SHAP values. For the Dead or Deceased class, the top three most important features were identified as Age, Regional_Node_Examined, & Regional_Node_Positive. For the Surviving or Alive class, the top three most important features are Age, Regional_Node_Examined, & Regional_Node_Positive (Figure: 11). We also examined the top three features of this dataset and their associated epistemic uncertainty, sign stability, and SHAP value distribution with mean SHAP and 95% SHAP confidence intervals. For the Surviving or Alive category, the epistemic uncertainties for the features Age, Regional_Node_Examined, & Regional_Node_Positive are 0.042, 0.044, and 0.037, respectively. The sign stabilities for these features are 98.3%, 94.7%, and 92.0%, respectively (Figure: 12). For the Deceased or Dead category, the epistemic uncertainties for the features Age, Regional_Node_Examined, &
Regional_Node_Positive are 0.042, 0.044, and 0.037, respectively (Figure  13). The sign stabilities for these features are 98.3%, 94.7%, and 92.0%, respectively. For both classes, the metric reflects model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. SHAP values are non-uniform for Age, Regional_Node_Examined, & Regional_Node_Positive features, signaling high uncertainty about the features’ impact, along with a relatively dispersed distribution. This reinforces the model’s uncertainty about precise attribution. The sign stabilities indicate high consistency with low epistemic uncertainty.
6 Conclusion
This research study decomposes SHAP into two categories of uncertainty quantification: aleatoric and epistemic. Breaking down uncertainty allowed us to determine whether it arises from uneliminatable noise in the data or from a lack of knowledge about the true data/model distribution. We also introduced an entanglement term that captures the interaction or covariance between data and model uncertainties. Our approach provides a deeper understanding of SHAP uncertainty in terms of intervals and enables investigation of its origin. In domains such as healthcare, where the consequences can be significant, our method is useful. The proposed approach captures uncertainties that simple intervals cannot while aligning with modern uncertainty quantification practices. DST quantifies both certainty ($Bel$) and possibility ($Pl$) for SHAP attributions, allowing for the explicit modeling of ignorance and epistemic uncertainty. This is particularly useful in ensemble bagging models with conflicting feature attributions. DST offers a non-probabilistic method of expressing confidence in SHAP values, which is useful when data is scarce or evidence is subjective. The framework extends beyond classical probability, offering tools for interpreting and managing uncertainty in model explanations. Identifying constrained SHAP intervals is often challenging in practice for large or complex models [86]. The proposed framework aims to facilitate the estimation of simple uncertainty intervals by leveraging the structural properties of belief functions, uncertainty theorems, and Dirichlet processes. These processes can be efficiently sampled or approximated, thereby enhancing the framework’s efficacy and accessibility. The framework provides control parameters, such as $\alpha$, which enable users to control the exploration-exploitation process of the hypothesis space. Users can choose balanced exploration, exploration of the best-performing trees, or uniform uncertainty estimation. Furthermore, uncertainty theory quantifies confidence in attribution magnitude, with entropy minimization guiding optimal data acquisition. This could facilitate explanations of SHAP values that account for uncertainty, even in high-dimensional or real-world settings. Additionally, our framework enhances uncertainty reporting. In a healthcare context, for example, SHAP values with 95% confidence intervals replace point estimates. The confidence-triggered verification is essential for determining which features require a domain expert’s review. Features with $phi_{i}<0.8$ require a review from a domain expert. This work also touches on the boundaries of AI regulations which emphasizes on the fact that the AI frameworks must provide auditable uncertainty metrics [87]. Together, variance, entropy, and sign stability provide a complete picture of uncertainty. Each chart provides insight into the stability of a SHAP attribution. The user can examine standard deviation, entropy, and sign stability. A mean $\pm 2\sigma$ SHAP indicates high epistemic uncertainty if SHAP varies widely across sub-models. The SHAP confidence interval represents the distribution of SHAP values; multimodality or flatness indicates disagreement across model variants. Uncertainty metrics quantify the reliability and stability of a feature’s attribution.
Acknowledgments
One of the authors of this work has been financially supported by the German Federal Ministry of Health (BMG) under grant No.: ZMI5- 2523GHP027 (project “Strengthening National Immunization Technical Advisory Groups and their Evidence-based Decision-making in the WHO European Region and Globally” SENSE) part of the Global Health Protection Programme, GHPP.

### v2

@@ ACKNOWLEDGMENTS
  • Figure 1: The SHAP summary plot provides a visual representation of the impact of various features (fifteen most contributing features) on a model’s prediction for the Short Stay, Medium Stay, & Long Stay class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The blue shaded area indicates the 95% confidence interval($\pm 2\sigma$) along with the horizontal dashed line of the feature’s impact on the prediction. The plot suggests that NumTransfers, NumNotes, NumCallouts are the most impactful feature, significantly contributing to the high probability of the prediction’s shift towards the Short Stay class; NumTransfers, NumCallouts, NumNotes are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Medium Stay class; NumTransfers, NumNotes, NumCallouts are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Long Stay class The features with mean SHAP value on the right side of the vertical dashed line contribute positively towards the prediction and vice-versa.
@@
  • Figure 3: The most contributing features to the Medium Stay class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
@@
  • Figure 5: This figure illustrates the relationship between a feature’s importance, measured by the mean absolute SHAP value, and the uncertainty surrounding its explanation, quantified by the standard deviation of SHAP values across a model ensemble, for various model prediction classes for the MIMIC-III Dataset. A high positive correlation ($\rho$) indicates that the model’s most significant features are also the most unstable, meaning their attributed contribution to the prediction varies significantly. Conversely, a low correlation suggests that feature importance and explanation stability are decoupled; the model can have important features it is certain about, and unimportant features it is uncertain about. This analysis is important for assessing the reliability of a model’s explanations, especially in high-stakes contexts.
@@
  • Figure 6: The SHAP summary plot provides a visual representation of the impact of the fifteen features with highest SHAP feature importance on a model’s prediction for the No Risk, Low Risk, Medium Risk, & High Risk class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented. The blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction. The plot suggests that Alcohol, SocioeconomicStatus, PreviousTreatment are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the No Risk class; SocioeconomicStatus, Age, Symptom are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Low Risk class; SocioeconomicStatus, Alcohol, PreviousTreatment are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Medium Risk class; SocioeconomicStatus, Alcohol, MenstrualHistory are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the High Risk class.
@@
  • Figure 11: This figure illustrates the relationship between a feature’s importance, measured by the mean absolute SHAP value, and the uncertainty surrounding its explanation, quantified by the standard deviation of SHAP values across a model ensemble, for various model prediction classes for the Ovarian Cancer Dataset. A high positive correlation ($\rho$) indicates that the model’s most significant features are also the most unstable, meaning their attributed contribution to the prediction varies significantly. Conversely, a low correlation suggests that feature importance and explanation stability are decoupled; the model can have important features it is certain about, and unimportant features it is uncertain about. This analysis is important for assessing the reliability of a model’s explanations, especially in high-stakes contexts.)
@@
  • Figure 12: The SHAP summary plot provides a visual representation of the impact of various features (fifteen most contributing) on a model’s prediction for the Deceased & Alive class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented. The blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction along with the horizontal dashed line of the feature’s impact on the prediction. The plot suggests that Age, Regional_Node_Examined, Tumor_Size are the most impactful features, significantly contributing to the high probability of the prediction’s shift towards both Deceased & Alive class, although in different directions. SHAP value on the right side of the vertical dashed line contribute positively towards the prediction and vice-versa.
@@
  • Figure 13: The distribution of SHAP values for the four most contributing features to the Alive class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
@@
  • Figure 15: This figure illustrates the relationship between a feature’s importance, measured by the mean absolute SHAP value, and the uncertainty surrounding its explanation, quantified by the standard deviation of SHAP values across a model ensemble, for various model prediction classes for the SEER Breast Cancer Dataset. A high positive correlation ($\rho$) indicates that the model’s most significant features are also the most unstable, meaning their attributed contribution to the prediction varies significantly. Conversely, a low correlation suggests that feature importance and explanation stability are decoupled; the model can have important features it is certain about, and unimportant features it is uncertain about. This analysis is important for assessing the reliability of a model’s explanations, especially in high-stakes contexts.